OpenStudy (anonymous):

Simplify the expression: ln10-ln(x+6)=5 I know that the answer is: {10-6e^5/e^5} But I don't know how to work out the problem.

OpenStudy (alfie):

Log's properties and exponential function. What happens when you apply a function and its inverse?

OpenStudy (anonymous):

I'm not sure exactly.

OpenStudy (jeyekomon):

Just use the basic rule for the logarithm: ln(a) - ln(b) = ln(a / b) ln10 - ln(x+6) = 5 ln(10 / (x+6)) = 5 Then use the inverse function to logarithm - the exponential function: ln(10 / (x+6)) = 5 exp( ln(10 / (x+6)) ) = exp(5) 10 / (x+6) = exp(5) 10 / exp(5) = x + 6 (10 - 6*exp(5)) / exp(5) = x

OpenStudy (jeyekomon):

Oh and by exp(5) I mean e^5 of course. It's the same thing.

OpenStudy (anonymous):

Thank you :)